# Hydrodynamic noise in one dimension: projected Kubo formula and its vanishing in integrable models

### v1

@@ 4 Constitutive relations and the hydrodynamic equation
In the construction of [10], the Onsager matrix $\mathfrak{L}_{o_{1},o_{2}}=(o_{1},o_{2})_{\rm dif}$ is written as an inner product in $\mathcal{H}_{\rm dif}$. There is a subspace, $\mathcal{H}_{\rm scat}$, that is spanned by the image of a certain bilinear map on the conserved quantities. This subspace is physically interpreted as describing the aggregated effect of nonlinear two-wave scattering events of hydrodynamic modes emanating from initial fluctuations and transported by the Euler equation. This nonlinear scattering and deterministic transport of fluctuations is the source of long-range correlations, as explained in [24]. As these have $\mathcal{O}(\ell^{-1})$ strength (see (22)), they contribute to $\mathfrak{L}_{o_{1},o_{2}}$ at the same order as the correlation between hydrodynamic noise $\xi_{o_{1}}$ and $\xi_{o_{2}}$. This is why this subspace must be subtracted to obtain the noise covariance.
Point III is expected – for current observables $o=j_{k}$, the diffusion functions $\hat{\mathfrak{D}}_{k}^{~{}i}(\underline{q}(x,t))$ (using the notation (26)) are usually assumed to be related to the noise covariance by Einstein’s relation. However, as far as we are aware, this has not been derived before in the context of (3.1). Further, here the formula is more general because it considers diffusion coefficients for arbitrary observables.
@@
We note that in [23], the same projected Kubo formula (40) is obtained for linearly degenerate systems by using a fluctuation principle of the type (3.1) but under the diffusive scaling, instead of the large-deviation scaling. Importantly, the result is verified by explicit calculations in a stochastic model, thus confirming the projected Kubo formula. The calculation methods are quite different, because of the different scaling.
4.1 The hydrodynamic equations at the diffusive scale
We first write down the diffusive-scale hydrodynamic equation from (3.1), taking into account the results above, Points I-III.
The hydrodynamic equation is obtained from this and the conservation laws, written for the long-wavelength state $\langle\cdots\rangle_{\ell}$ (see (13)), whose corresponding average in the emergent fluctuating theory is denoted $\langle\!\langle\cdots\rangle\!\rangle_{\ell}$:
Inserting (3.1) for the case $o=j_{i}$ is this, using $\langle\!\langle\xi_{o}(x,t)\rangle\!\rangle_{\ell}=0$ and the fact that for the diffusive order $\ell^{-1}$ we may use the leading cumulant-expansion form
$\partial_{t}\mathsf{q}_{i}(x,t)+\partial_{x}\langle\!\langle\mathsf{j}_{i}^{\rm reg}(\underline{q}(x,t))\rangle\!\rangle_{\ell}=\frac{1}{2\ell}\sum_{k}\hat{\mathfrak{D}}_{i}^{~{}k}(\underline{\mathsf{q}}(x,t))\partial_{x}\mathsf{q}_{k}(x,t)$ (47)
$\partial_{t}\mathsf{q}_{i}(x,t)+\sum_{k}\mathsf{A}_{i}^{~{}k}(x,t)\partial_{x}\mathsf{q}_{k}(x,t)$ (48)
$+\,\frac{1}{4\ell}\sum_{r,r^{\prime},s,s^{\prime},\pm}\partial_{x}\Big{(}\langle j_{i}^{-},Q_{r},Q_{s}\rangle^{\rm c}_{\underline{\beta}(x,t)}\mathsf{C}^{rr^{\prime}}(x,t)\mathsf{C}^{ss^{\prime}}(x,t)E_{r^{\prime}s^{\prime}}(x\pm 0^{+},x\mp 0^{+};t)\Big{)}$
$=\ \frac{1}{2\ell}\sum_{k}\hat{\mathfrak{D}}_{i}^{~{}k}(\underline{\mathsf{q}}(x,t))\partial_{x}\mathsf{q}_{k}(x,t)$
with
$\partial_{t}E_{rs}(x,y;t)+\partial_{x}\big{(}\mathsf{A}_{r}^{~{}k}(x,t)E_{ks}(x,y;t)\big{)}+\partial_{y}\big{(}E_{rk}(x,y;t)\mathsf{A}_{s}^{~{}k}(y,t)\big{)}$ (49)
$\qquad\qquad\qquad\qquad=-\delta(x-y)\sum_{kl}\langle j_{k},Q_{r},Q_{s}\rangle^{\rm c}_{\underline{\beta}(x,t)}\mathsf{C}^{kl}(x,t)\partial_{x}\mathsf{q}_{l}(x,t).$
Here $\beta^{i}(x,t):=\beta^{i}(\underline{\mathsf{q}}(x,t))$ (see Eq. (12)) and likewise for the hydrodynamic matrices $\mathsf{A}_{i}^{~{}k}(x,t)$ and $\mathsf{C}_{ik}(x,t)$. In the state (13), the initial conditions are
$\beta^{i}(x,0)=\beta^{i}(x),\quad E_{rs}(x,y;0)=0$ (50)
but these equations are more general, and valid also for initial states that admit non-trivial long-range correlations $E_{rs}(x,y;0)\neq 0$. Eqs. (48), (49) generalise [19, Eqs 9,10], which was for integrable models, to the inclusion of the bare diffusion generically present away from integrability.
@@ 4.2 Correction to the microcanonical average and point splitting
Let us consider the microcanonical correction $\delta\mathsf{o}({\underline{q}}(x,t))$, Eq. (30), for one-point averages generically in non-equilibrium states. It be evaluated by studying the average of $o=o(0,0)$ in a stationary, maximal entropy state $\langle\cdots\rangle_{\underline{\beta}}$, and taking into account that it has short-range correlations. Recall that by the equivalence of ensembles,
@@
[…]
@@
Consider a more general non-equilibrium state $\langle\cdots\rangle_{\ell}$ such as (13), and recall the property (37). Then a similar calculation as (53) and (54) shows that, in every such state, $L^{-1}\delta\mathsf{o}\big{(}{\underline{q}}\big{)}$ cancels the $1/L$ singularities of $\mathsf{o}(\underline{q}(x,t))$. But we can implement this cancellation simply as
@@
As we can reproduce correlation functions by functional differentiations with respect to external fields in such states (see [42, 26]), this regularisation works for correlation functions as well. This is a universal procedure: the explicit $L$ dependence has disappeared, and all equal-point singularities are avoided.
@@ 4.3 Noise correlations via projected Kubo formula
We now evaluate the correlation matrix $\hat{\mathfrak{L}}_{o,o^{\prime}}$ in (20). We will show that, in one dimension, it is not the usual Kubo formula, but instead a projected Kubo formula, where the quadratic charge subspace [10] has been projected out.
[…]
@@
$=$ $\ell\int_{-\infty}^{\infty}{\rm d}t\,\Big{(}\ell\int{\rm d}x\,\langle\!\langle o_{1}(x,t),o_{2}(0,0)\rangle\!\rangle^{\rm c}-\mathsf{D}_{o_{1},o_{2}}\Big{)}$
where we have written it in terms of the emergent fluctuating theory, with scaled space-time coordinates, where we will use (3.1). We note that the Drude weight (42) can be written as
@@
Using (3.1), we analyse the terms that contribute to $\langle\!\langle o_{1}(x_{1},t_{1}),o_{2}(x_{2},t_{2})\rangle\!\rangle^{\rm c}$ at orders $\ell^{-1}$ and $\ell^{-2}$; and in particular, those that contribute to (60) (equivalently (62)), hence that are nonzero under space-time integration. For each of the observables $o_{1}$ and $o_{2}$, we consider the microcanonical, diffusive and noise terms, and evaluate their correlations.
@@
$\mathfrak{L}_{o_{1},o_{2}}$ =$\stackrel{{\scriptstyle\text{micro-micro}}}{{=}}$ $\frac{1}{2}\sum_{I,I^{\prime}}\int{\rm d}t\int{\rm d}x\,\langle o_{1}^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}\langle o_{2}^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}\delta(x-\mathsf{v}_{I}t)\delta(x-\mathsf{v}_{I^{\prime}}t)$ (69)
@@ 4.4 Extended Einstein relation
Finally, we establish the relation between the diffusion function $\hat{\mathfrak{D}}_{o}^{~{}i}(\underline{q})$ in (3.1) (equiv. (3.1)), and the noise covariance $\hat{\mathfrak{L}}_{o,j_{i}}(\underline{q})$ for the observable $o$ and currents $j_{i}$. For this purpose, let us write, using space-time translation invariance of the state and the conservation laws,
@@ 5 Absence of noise in integrable systems
It was recently conjectured that the noise in integrable systems must vanish, see [20, 21, 19]. In particular, in [19] a nonlinear Boltzmann-Gibbs principle in the ballistic $1/\ell$ expansion, as in (3.1), is used without noise, in order to evaluate diffusive-order corrections to the hydrodynamic equation, showing agreement with numerical simulation. That is, in intregable models, the contribution to the diffusive scale comes solely from the long-range correlations induced by the large-scale variations of the initial state [24], which are of order $1/\ell$ and affect the (point-splitted) microcanonical average.
In this section we provide a proof that this is the case. First, we show this using previous results on the Onsager matrix, and then we provide a first-principle argument, based on the fluid-cell relaxation picture of Sec. 3.2. We show that in integrable systems, the noise $\xi_{j_{i}}$ for the currents $j_{i}$, as appears at order $\ell^{-1}$ in (3.1), vanishes
@@
that is,
$\hat{\mathfrak{L}}_{o,j_{i}}(\underline{\mathsf{q}})=\hat{\mathfrak{L}}_{j_{i},o}(\underline{\mathsf{q}})=0,\quad\hat{\mathfrak{D}}_{o}^{~{}k}(\underline{\mathsf{q}})=0$ (91)
where the second relation follows from (43).
First proof. We show that this is a consequence of our main result (40), along with the following observation.
In the hard rods model, the Onsager matrix for currents $\mathfrak{L}_{j_{i},j_{k}}$, which is in fact an operator as it is taken in the basis of the conserved hard rods velocity $i=v\in{\mathbb{R}}$, was evaluated exactly, see e.g. [28, 4]. Further, in [29, 30] (see also [31]) the Onsager matrix for currents $\mathfrak{L}_{j_{i},j_{k}}$ was evaluated in integrable models, generalising the hard-rod formula. This was done using a variety of techniques, including a direct calculation from form factors in [30]. Then, it was shown in [10] that the general formula in fact takes the form
$\mathfrak{L}_{j_{i},j_{k}}\stackrel{{\scriptstyle\text{integrable models}}}{{=}}\frac{1}{2}\sum_{I\neq I^{\prime}}\frac{\langle j_{i}^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}\langle j_{k}^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}}{|\mathsf{v}_{I}-\mathsf{v}_{I^{\prime}}|}.$ (92)
@@
$(\hat{\mathfrak{L}}_{o,j_{i}})^{2}\leq\hat{\mathfrak{L}}_{o,o}\hat{\mathfrak{L}}_{j_{i},j_{i}}\stackrel{{\scriptstyle\text{integrable models}}}{{=}}0$ (94)
@@
Second proof. We provide a general first-principle argument based on relaxation of fluid-cell means and fundamental properties of integrable systems.
In many-body integrable models, there exists an infinite number of commuting, independent “higher” time flows, besides the “usual” time evolution. If the integrable system is Hamitonian, then these are generated by infinitely many linearly and functionally independent commuting, extensive Hamiltonians. In translation-invariant systems, these time flows are also translation invariant. That is, we can evolve any local observable not just in time $t=t_{1}$, but also in times $t_{2},t_{3},\ldots$, which we denote by the commuting maps $\tau_{2}^{t_{2}},\tau_{3}^{t_{3}},\ldots$:
$o(x,t_{1},t_{2},t_{3},\ldots)=\tau_{2}^{t_{2}}\tau_{3}^{t_{3}}\cdots o(x,t_{1}).$ (95)
@@
$\partial_{t_{k}}q_{i}(x,t_{1},\ldots,t_{n})+\partial_{x}j_{ki}(x,t_{1},\ldots,t_{n})=0.$ (96)
Further, using this, it was shown in [10, App 4] that the Onsager matrix is invariant under these flows,
$\mathfrak{L}_{\tau_{2}^{t_{2}}\tau_{3}^{t_{3}}\cdots j_{i},o}=\mathfrak{L}_{j_{i},o}.$ (97)
@@
$\hat{\mathfrak{L}}_{\tau_{2}^{t_{2}}\tau_{3}^{t_{3}}\cdots j_{i},o}=\hat{\mathfrak{L}}_{j_{i},o}.$ (98)
Now recall the argument in Sec. 3.2 for the origin of hydrodynamic noise. On a fluid cell of size $L$, after a time $T$, there has been $\mathcal{O}(LT)=\mathcal{O}(\ell)$ fluctuation events for the fluid-cell mean $\overline{o}$ of the observable $o$, while, thanks to the conservation laws, the conserved densities $\overline{q_{i}}$’s are essentially constant. Hence, taking an average over a time $t\in[0,\tau]$ (or, equivalently, not taking the time average, but using typicality) relaxation to the microcanonical average occurs, plus a noise of variance $\propto 1/\ell$ by the central limit theorem.
Consider $o=j_{i}$ to be a current. Let us change this observable by performing an averaging over $n$ of the higher time flows, with higher times in the range $[0,T^{\prime}]$ for some fixed $T^{\prime}>1$ (not scaling with $\ell$),
$[j_{i}]=\frac{1}{{T^{\prime}}^{n}}\int_{0}^{T^{\prime}}{\rm d}t_{2}\cdots\int_{0}^{T^{\prime}}{\rm d}t_{n+1}\prod_{k=2}^{n+1}\tau_{k}^{t_{k}}j_{i}.$ (99)
Taking the time-average over $t_{1}\in[0,T]$, the fluid-cell mean $\overline{[j_{i}]}$ is the average of the original fluid cell mean $\overline{j_{i}}$ over the time-hyperbox $[0,T]\times[0,T^{\prime}]^{n}$ of dimension $n+1$, with $t_{1}\in[0,T]$ and $t_{k}\in[0,T^{\prime}]$ for $k=2,\ldots,n+1$. On this hyperbox, fluid cell means of conserved densities $\overline{q_{i}}$’s are all essentially constant, because of (96). As each time flow is independent, the fluctuations induced, from typical configurations, by each flow are also independent. Thus, on the space-time hyperbox $[0,L]\times[0,T]\times[0,T^{\prime}]^{n}$ the configuration displays a number of independent fluctuation events (particle collisions, etc.) proportional to the volume of this box, hence $\mathcal{O}(LT{T^{\prime}}^{n})$. That is, the fluid cell mean $\overline{[j_{i}]}$ has undergone ${T^{\prime}}^{n}$ times more fluctuations than $\overline{j_{i}}$. Taking $n$ large after the $\ell\to\infty$ asymptotic, we thus find by the central limit theorem that the noise for the observable $[j_{i}]$ has variance
$\langle\!\langle\xi_{[j_{i}]}(x,t)\xi_{[j_{i}]}(x^{\prime},t^{\prime})\rangle\!\rangle\sim c(T^{\prime})\ell^{-1}\delta(x-x^{\prime})\delta(t-t^{\prime}),\quad c(T^{\prime})=\mathcal{O}({T^{\prime}}^{-n})$ (100)

### v2

@@ 4 Constitutive relations
In [24], the subspace $\mathcal{H}_{\rm scat}$ is physically interpreted as describing the aggregated effect of nonlinear two-wave scattering events of hydrodynamic modes emanating from initial fluctuations and transported by the Euler equation. Nonlinear scattering of deterministically transported fluctuations is the source of long-range correlations, as explained in [7]. As these have $\mathcal{O}(\ell^{-1})$ strength (see (20)), they contribute to $\mathfrak{L}_{o_{1},o_{2}}$ at the same order as the correlation between hydrodynamic noise $\xi_{o_{1}}$ and $\xi_{o_{2}}$. This is why this subspace must be subtracted to obtain the noise covariance. Thus, the present result finally puts together these various concepts in a coherent theory.
Point III is expected – for current observables $o=j_{k}$, the diffusion functions $\hat{\mathfrak{D}}_{k}^{~i}(\underline{q}(x,t))$ (using the notation (27)) are usually assumed to be related to the noise covariance by the Einstein relation. However, as far as we are aware, this has not been derived before in the context of (21).
@@
We note that in [39], the same projected Kubo formula (44) is obtained for linearly degenerate systems by using a fluctuation principle of the type (21) but under the diffusive scaling, instead of the large-deviation scaling. Importantly, the result is verified by explicit calculations in a stochastic model, thus confirming the projected Kubo formula. The calculation methods are quite different, because of the different scaling.
@@ 4.1 Correction to the microcanonical average and point splitting
Let us consider the microcanonical correction $\delta\mathsf{o}({\underline{q}}(x,t))$, Eq. (36), for one-point averages generically in non-equilibrium states. It can be evaluated by studying the average of $o=o(0,0)$ in a stationary, maximal entropy state $\langle\cdots\rangle_{\underline{\beta}}$, and taking into account that it has short-range correlations. Recall that by the equivalence of ensembles,
@@
where we used homogeneity to write the average as an average of a fluid-cell mean, and then used (21) and the vanishing of noise and diffusive parts in a maximal entropy state. We find that the correction term should simply “allow” for the averaging to pass inside the function $\mathsf{o}(\underline{q})$. Recall that $\langle\!\langle\cdots\rangle\!\rangle$ implicitly depends on $L$ because the fluctuating fields $q_{i}(x,t)$ represents fluctuating fluid-cell means. Let us evaluate equal-point connected correlation functions of these fluctuating fields using their fluid-cell mean definition. Assuming short-range correlations,
@@
Consider a more general non-equilibrium state $\langle\cdots\rangle_{\ell}$ such as (12), and recall the property (41). Then a similar calculation as (54) and (55) shows that, in every such state, $L^{-1}\delta\mathsf{o}\big{(}{\underline{q}}\big{)}$ cancels the $1/L$ singularities of $\mathsf{o}(\underline{q}(x,t))$. In this calculation, Eq. (41) is used only to evaluate the leading $1/L$ correction, hence we can neglect the $1/\ell$ correction in the delta-function coefficient in (41). But we can implement this cancellation simply as
@@
Finally, as we can reproduce correlation functions by functional differentiations with respect to external fields in such states (see [55, 43]), this regularisation works for correlation functions as well. This is a universal procedure: the explicit $L$ dependence has disappeared, and all equal-point singularities are avoided.
@@ 4.2 Noise correlations via projected Kubo formula
We now evaluate the correlation matrix $\hat{\mathfrak{L}}_{o,o^{\prime}}$ in (32). We will show that it is not the usual Kubo formula, but instead a projected Kubo formula, where the quadratic charge subspace [24] has been projected out.
For this purpose, let us consider again a stationary, maximal entropy state $\langle\cdots\rangle$, and evaluate, using (21), the Kubo formula
@@
$=$ $\ell\int_{-\infty}^{\infty}{\rm d}t\,\Big{(}\ell\int{\rm d}x\,\langle\!\langle o_{1}(x,t),o_{2}(0,0)\rangle\!\rangle^{\rm c}-\mathsf{D}_{o_{1},o_{2}}\Big{)}.$
In the second line we have written it in terms of the emergent fluctuating theory, with scaled space-time coordinates, where we will use (21). We note that the Drude weight (46) can be written as
@@
Using (21), we analyse the terms that contribute to $\langle\!\langle o_{1}(x_{1},t_{1}),o_{2}(x_{2},t_{2})\rangle\!\rangle^{\rm c}$ at orders $\ell^{-1}$ and $\ell^{-2}$; and in particular, those that contribute to (61) (equivalently (63)), hence that are nonzero under space-time integration. For each of the observables $o_{1}$ and $o_{2}$, we consider the microcanonical, diffusive and noise terms, and evaluate their correlations.
@@
$\mathfrak{L}_{o_{1},o_{2}}$ $\stackrel{{\scriptstyle\text{micro-micro}}}{{=}}$ $\frac{1}{2}\sum_{I,I^{\prime}}\int{\rm d}t\int{\rm d}x\,\langle o_{1}^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}\langle o_{2}^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}\delta(x-\mathsf{v}_{I}t)\delta(x-\mathsf{v}_{I^{\prime}}t)$ (70)
@@ 4.3 Extended Einstein relation
Finally, we establish the relation between the diffusion function $\hat{\mathfrak{D}}_{o}^{~i}(\underline{q})$ in (21) (equiv. (21)), and the noise covariance $\hat{\mathfrak{L}}_{o,j_{i}}(\underline{q})$ for the observable $o$ and currents $j_{i}$. For this purpose, let us write, using space-time translation invariance of the state and the conservation laws,
@@ 5 Absence of noise in integrable systems
It was recently conjectured that the noise in integrable systems must vanish, see [36, 37, 25]. In particular, in [25] a nonlinear Boltzmann-Gibbs principle in the ballistic $1/\ell$ expansion, as in (21), is used without noise, in order to evaluate diffusive-order corrections to the hydrodynamic equation, showing agreement with numerical simulation. That is, in intregable models, the contribution to the diffusive scale comes solely from the long-range correlations induced by the large-scale variations of the initial state [7], which are of order $1/\ell$ and affect the (point-splitted) microcanonical average.
In this section we provide a proof that this is the case. We show that, under an appropriate choice of the current observables in integrable models – a choice of “gauge” –, which we denote $j_{i}]_{\infty}$, the nonlinear fluctuating Boltzmann-Gibbs principle simplifies to a non-fluctuating one at all orders of the hydrodynamic expansion:
[…]
At the diffusive order, with $\mathcal{O}(\ell^{-2})$ instead of $\mathcal{O}(\ell^{-\infty})$, the result holds for the PT-symmetric current $j_{i}$ instead of $j_{i}]_{\infty}$ (that is, without the need to further fix the gauge).
[…]
First, we show the result at the diffusive order (92), with the PT-symmetric current $j_{i}$. This is done using previous, independent results on the Onsager matrix. We show that in integrable systems, the noise $\xi_{j_{i}}$ for the currents $j_{i}$, as appears at order $\ell^{-1}$ in (21), vanishes, and as a consequence, the bare diffusion for all observables also vanishes. That is,
@@
and more specifically
$\hat{\mathfrak{L}}_{j_{i},j_{i}}(\underline{\mathsf{q}})=0\ \Rightarrow\ \hat{\mathfrak{L}}_{o,j_{i}}(\underline{\mathsf{q}})=\hat{\mathfrak{L}}_{j_{i},o}(\underline{\mathsf{q}})=0\ \Rightarrow\ \hat{\mathfrak{D}}_{o}^{~k}(\underline{\mathsf{q}})=0$ (94)
where the last equality follows from (47).
Second, we give a first-principle argument, based on the fluid-cell relaxation picture of Sec. 3.2. This argument is important in two ways:
(1) It provides an independent, first-principle derivation of the result for the Onsager matrix in integrable models, Eq. (95). Indeed, as we have shown that the noise strength is the projected Kubo formula, a corollary that this vanishes is that the un-projected, full Kubo formula is given by its projection onto quadratic charges. This reproduces the known expression for the Onsager matrix in integrable systems in terms of quasi-particles.
(2) It allows us to show that in fact, the noise for currents must vanish at all orders in the hydrodynamic expansion, Eq. (91). This is true, beyond the diffusive scale, as long as current observables are chosen appropriately. Recall that charge densities and currents are defined only up to total derivatives of local observables – this is the “gauge” freedom in determining densities and currents. At the diffusive level, this gauge freedom is lifted by choosing the PT gauge [41]. But beyond it, there is no natural gauge, see however a discussion in [64]. Our proof provides a partially constructive way of choosing the gauge, that guarantees that the nonlinear Boltzmann-Gibbs principle for currents does not admit noise at all orders of the hydrodynamic expansion.
First proof. We show that the absence of noise at the diffusive scale, Eq. (92) or equivalently (94), is a consequence of our main result (44), along with the following observation.
In the hard rods model, the Onsager matrix $\mathfrak{L}_{j_{i},j_{k}}$, which is in fact an integral operator as it is taken in the basis of the conserved hard rods velocity $i=v\in{\mathbb{R}}$, was evaluated exactly, see e.g. [45, 5, 65]. Further, in [40, 41] the Onsager matrix $\mathfrak{L}_{j_{i},j_{k}}$ was evaluated in integrable models using quantum form factor techniques, generalising the hard-rod formula (see also [42] where the diagonal part was evaluated using a quasi-particle picture). Then, it was shown in [24] that this general formula in fact corresponds to a projection, within the diffusive Hilbert space $\mathcal{H}_{\rm dif}$, onto the scatteing space formed by quadratic charges (see the discussion in Sec. 4)
$\mathfrak{L}_{j_{i},j_{k}}\stackrel{{\scriptstyle\text{integrable models}}}{{=}}\big{(}\mathbb{P}_{\rm scat}j_{i},\mathbb{P}_{\rm scat}j_{k}\big{)}_{\rm dif}=\frac{1}{2}\sum_{I\neq I^{\prime}}\frac{\langle j_{i}^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}\langle j_{k}^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}}{|\mathsf{v}_{I}-\mathsf{v}_{I^{\prime}}|}.$ (95)
@@
[…]
@@
Second proof. We show (92), and then more generally (91), using a general first-principle argument based on relaxation of fluid-cell means and fundamental properties of integrable systems.
In many-body integrable models, there exists an infinite number of commuting, independent “higher” time flows, besides the “usual” time evolution. If the integrable system is Hamitonian, then these are generated by infinitely many linearly and functionally independent commuting, extensive Hamiltonians. In translation-invariant systems, these time flows are also translation invariant. That is, we can evolve any local observable not just in time $\mathrm{t}=\mathrm{t}_{1}$, but also in times $\mathrm{t}_{2},\,\mathrm{t}_{3},\,\ldots$, which we denote by the commuting one-parameter groups of maps $\tau_{2}^{\mathrm{t}_{2}},\tau_{3}^{\mathrm{t}_{3}},\ldots$ acting on observables:
$o(\mathrm{x},\mathrm{t}_{1},\mathrm{t}_{2},\mathrm{t}_{3},\ldots)=\tau_{2}^{\mathrm{t}_{2}}\tau_{3}^{\mathrm{t}_{3}}\cdots o(\mathrm{x},\mathrm{t}_{1}),\quad\mathrm{t}_{1},\mathrm{t}_{2},\mathrm{t}_{3},\ldots\in{\mathbb{R}}.$ (98)
@@
$\partial_{\mathrm{t}_{k}}q_{i}(\mathrm{x},\mathrm{t}_{1},\ldots,t_{n})+\partial_{\mathrm{x}}j_{ki}(\mathrm{x},\mathrm{t}_{1},\ldots,\mathrm{t}_{n})=0.$ (99)
Consistency requires $\partial_{\mathrm{t}_{k}}j_{li}=\partial_{\mathrm{t}_{l}}j_{ki}$. Further, using this, it was shown in [24, App 4] that the Onsager matrix is invariant under these flows,
$\mathfrak{L}_{\tau_{2}^{\mathrm{t}_{2}}\tau_{3}^{\mathrm{t}_{3}}\cdots j_{i},o}=\mathfrak{L}_{j_{i},o}.$ (100)
@@
$\hat{\mathfrak{L}}_{\tau_{2}^{\mathrm{t}_{2}}\tau_{3}^{\mathrm{t}_{3}}\cdots j_{i},o}=\hat{\mathfrak{L}}_{j_{i},o}.$ (101)
Now recall the argument in Sec. 3.2 for the origin of hydrodynamic noise. On a fluid cell of size $L$, after a time $T$, there has been $\mathcal{O}(LT)=\mathcal{O}(\ell)$ fluctuation events for the fluid-cell mean $\overline{o}$ of the observable $o$, while, thanks to the conservation laws, the conserved densities $\overline{q_{i}}$’s are essentially constant; see Fig. 2. Hence, taking an average over a time $t\in[0,T]$ (or, equivalently, not taking the time average, but using typicality) relaxation to the microcanonical average occurs, plus a noise of variance $\propto 1/\ell$ by the central limit theorem.
Consider $o=j_{i}$ to be a current. Let us change this observable by averaging over $n$ of the higher time flows, with higher times in the range $0,T^{\prime}]$ for some fixed $T^{\prime}>1$ (not scaling with $\ell$),
$j_{i}]=\frac{1}{{T^{\prime}}^{n}}\int_{0}^{T^{\prime}}{\rm d}\mathrm{t}_{2}\cdots\int_{0}^{T^{\prime}}{\rm d}\mathrm{t}_{n+1}\prod_{k=2}^{n+1}\tau_{k}^{\mathrm{t}_{k}}j_{i}.$ (102)
Taking the time-average over $\mathrm{t}_{1}\in[0,T]$, the fluid-cell mean $\overline{[j_{i}]}$ is the average of the original fluid cell mean $\overline{j_{i}}$ over the time-hyperbox $0,T]\times[0,T^{\prime}]^{n}$ of dimension $n+1$, with $\mathrm{t}_{1}\in[0,T]$ and $\mathrm{t}_{k}\in[0,T^{\prime}]$ for $k=2,\ldots,n+1$. On this hyperbox, fluid cell means of conserved densities $\overline{q_{i}}$’s are all essentially constant, because of (99). As each time flow is independent, the fluctuations induced, from typical configurations, by each flow are also independent. Thus, on the space-time hyperbox $0,L]\times[0,T]\times[0,T^{\prime}]^{n}$ the configuration shows a number of independent fluctuation events (particle collisions, etc.). In the higher-dimensional equivalent of Fig. 2, these fluctuation events lie on hypersurfaces of dimension $n+1$ criss-crossing the hyperbox. There are $\mathcal{O}(LT)$ such hypersurfaces, each of hypervolume $\propto{T^{\prime}}^{n}$. It is natural to assume that on each such hypervolume, there is typically ${T^{\prime}}^{n}/\ell_{\rm micro}^{n}$ independent changes, by an extension of the assumption of molecular chaos to this context with higher-dimensional time. Therefore, the fluid cell mean $\overline{[j_{i}]}$ has undergone $\propto{T^{\prime}}^{n}$ times more fluctuations than $\overline{j_{i}}$.
Taking $n$ and/or $T^{\prime}$ large after the $\ell\to\infty$ asymptotic, we thus find by the central limit theorem that the noise for the observable $j_{i}]$ has variance
$\langle\!\langle\xi_{[j_{i}]}(x,t)\xi_{[j_{i}]}(x^{\prime},t^{\prime})\rangle\!\rangle_{\ell}\sim c(T^{\prime})\ell^{-1}\delta(x-x^{\prime})\delta(t-t^{\prime}),\quad c(T^{\prime})=\mathcal{O}({T^{\prime}}^{-n})$ (103)